OpenStudy (anonymous):

Need a little bit of guidance on how to resolve this: A rope suspended from a ceiling supports an object of weight W at its opposite end. Another rope tied to the first at the middle is pulled horizontally with a force of 30N. The junction P of the ropes is in equilibrium. Calculate the weight W and the tension in the first rope.

OpenStudy (anonymous):

f=mg g= gravity

OpenStudy (anonymous):

@Jesther I know that. The difficulty I have here is the fact that only one variable was given (horizontal axis) and no angle. That makes it three unknowns. Really confusing...

OpenStudy (anonymous):

we already tackled that question, i forgot

OpenStudy (anonymous):

how do you mean, @Jesther ?

OpenStudy (anonymous):

what??

OpenStudy (anonymous):

you tackled this question, where?

OpenStudy (anonymous):

in our class

OpenStudy (anonymous):

And you forgot. That's really helpful.

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

is that net force??

OpenStudy (anonymous):

tension in the suspended rope and the weight of the object are what we need here.

OpenStudy (anonymous):

A rope suspended from a ceiling supports an object of weight W at its opposite end i think that is the clue word.. "opposite end"

OpenStudy (anonymous):

Does the following formula make sense? \[T=\sqrt{W^2+30^2}\]

OpenStudy (anonymous):

Then, weight in the rope would be: \[W_R=\frac{ W }{ \sqrt{1+(30/W)^2}}\]and force in the rope:\[F_R=\frac{ 30 }{ \sqrt{1+(W/30)^2} }\]\[T=F_R+W_R=\frac{ 30^2 }{ \sqrt{30^2+W^2} }+\frac{ W^2 }{ \sqrt{30^2+W^2}}=\sqrt{30^2+W^2}\]

OpenStudy (anonymous):

Thank you @CarlosGp, but that doesn't seem like a solution. We still have two unknowns.

OpenStudy (08surya):

pls give the diagram